Run Planning Meeting Notes, May 8 - May 14, 2025

From GlueXWiki
Jump to: navigation, search

The zoom link for the Hall D 0845 meeting is the first in this list

MCC whiteboard

Hall D counting room whiteboard

Hall D Logbook

Spreadsheet with list of runs

Back to Run Coordination Meetings:Spring2025#Runplan


Wednesday May 14

Last 24h

  1. No beam due to clogged magnet in arc 10
  2. Shift crew summaries
    • Day, 08:00-16:00 cancelled
    • Evening, 16:00-00:00 cancelled
    • Owl, 00:00-08:00 cancelled

Reports

  1. Follow-up Re: FDC wire efficiency at 8/9V thresholds Follow-up Re: Follow-up Re: FDC wire efficiency at 8/9V thresholds (Alex, Lubomir)
  2. Numbers of triggers collected since start of production running
    • PARA 0/90 deg: 1.84B
    • PERP 0/90 deg: 2.46B
    • PARA 45/135 deg: 2.49B
    • PERP 45/135 deg: 2.43B
    • AMO: 1.24B

Problems

  1. Updated VTP firmware (Sasha)

Plans

  1. Arc 10 magnet will be flushed starting 7:00AM
  2. Accelerator will change settings of helicity board between 10:00AM and 12:00PM (noon)
  3. Shifts will resume starting at 12:00PM (noon)
  4. Fill liquid-hydrogen target
  5. Hall-D Safety Walkthrough at 1:00PM
  6. When beam is back
    • Perform harp scan(s)
      • Try to improve beam convergence
      • In [1], Jay Benesch provided possible steps to achieve better convergence.
    • Check and tune trigger using AMO radiator
    • Continue production running
      • Start new sequence

Tuesday May 13

Last 24h

  1. No beam
    • Magnet in Arc 10 is overheating due to clogging
  2. Shift crew summaries
    • Day, 08:00-16:00 standby
    • Evening, 16:00-00:00 cancelled
    • Owl, 00:00-08:00 cancelled

Reports

  1. FDC wire efficiency at 8/9V thresholds (Lubomir)
  2. Low PS rates since last Wednesday (Alexandre)
  3. Numbers of triggers collected since start of production running
    • PARA 0/90 deg: 1.84B
    • PERP 0/90 deg: 2.46B
    • PARA 45/135 deg: 2.49B
    • PERP 45/135 deg: 2.43B
    • AMO: 1.24B

Problems

  1. Problems booting roccdc2 (Sergey)
  2. ECAL FCAL studies, Missing fADCs for high energy ECAL hits, Follow-up Re: ECAL FCAL studies, ECAL mapping (Beni, Justin, Mark)

Plans

  1. Shifts until Wednesday 12:00PM (noon) are cancelled
  2. Arc 10 magnet will be flushed Wednesday starting 7:00AM

Monday May 12

Last 24h

  1. No beam
    • Magnet in Arc 10 is overheating due to clogging
  2. Shift crew summaries

Reports

  1. CDC voltages restored (Beni)
  2. Target was switched off instead of just emptied
  3. Numbers of triggers collected since start of production running
    • PARA 0/90 deg: 1.84B
    • PERP 0/90 deg: 2.46B
    • PARA 45/135 deg: 2.49B
    • PERP 45/135 deg: 2.43B
    • AMO: 1.24B

Problems

  1. None

Plans

  1. Waiting for beam

Sunday May 11

Last 24h

  1. No beam
    • Repair of dipole power supply for our beam line [2]
  2. Shift crew summaries

Reports

  1. Online calibrations running (Sean)
    • Since 23:45 TRD is included in "PHYSICS_DIRC" DAQ configuration
  2. Follow-up Re: Confusion about BPM AD00 x,y position (tagger dump) in shift worker checklist (Naomi)
  3. Numbers of triggers collected since start of production running
    • PARA 0/90 deg: 1.84B
    • PERP 0/90 deg: 2.46B
    • PARA 45/135 deg: 2.49B
    • PERP 45/135 deg: 2.43B
    • AMO: 1.24B

Problems

  1. Rootspy issue: software is down? And cannot access Gluon159
    • Resolved by Sergey by rebooting gluon159.
  2. HallD Run #131492: CDC Alarm not clearing, CDC voltages restored, HallD Run #131492: CDC alarm PV:CDC:hv:G13 (Beni)

Plans

  1. Continue production running

Saturday May 10

Last 24h

  1. Accesses to Hall D and tagger vault to replace ADC card in CDC and repair amorphous motor
  2. 12:30 tuning of beam convergence
    • Since the beam optics for the whole accelerator was redone the convergence was far off
    • We worked with MCC for 6h and performed several harp scans: [3] [4] [5] [6] [7] [8] [9]
    • Final harp scan: [10] [11] [12] [13]
  3. 18:30 start data taking
    • Run 131481, raw mode with AMO radiator
    • 19:45 Resumed production data taking
    • Took 156M events for AMO and 294M events for PARA 0/90 deg. Then beam went away at 2:30 and did not return.
    • PARA 0/90 deg runs 131490, 131491, 131492 taken with EXPERT config that includes TRD [15]
  4. Shift crew summaries

Reports

  1. Successful replacement of ADC card in CDC Follow-up Re: HallD Run #131473: CDC occupancy plot, fADC125 replacment in ROCCDC2, Roc 26 (CDC2) multiple readout errors, could we try re-download of the daq config or reboot the crate?, [16] [17], [18] (Beni/Naomi)
  2. Missing BCAL LED front panel triggers (Zisis)
  3. FDC discr. thrsholds set to 8V, FDC discr. thresholds increased to 9V (Lubomir)
  4. Successful repair of stuck motor for amorphous radiators and harp wires Tagger amorphous radiator service Amorphous radiator stick problem (Hovanes)
  5. HallD Run #131483: Turn on Random Pulser
  6. Numbers of triggers collected since start of production running
    • PARA 0/90 deg: 1.84B
    • PERP 0/90 deg: 2.46B
    • PARA 45/135 deg: 2.49B
    • PERP 45/135 deg: 2.43B
    • AMO: 1.24B

Problems

  1. trig_client program not running?, Follow-up Re: trig_client program not running?

Plans

  1. Perform harp scan when beam back.
  2. Continue production running
    • Start with remaining 506M events for PARA 0/90 deg

Friday May 9

Last 24h

  1. No beam since Wed, May 7, 06:30
  2. MCC worked on beam-optics restore
  3. Beam back at about 2:30am
  4. Harp scan 1: [19][20][21][22]
    • Convergence not good
  5. Hard scan 2: [23][24][25]
    • Inconclusive because data from our harp missing
  6. No useful data taken
  7. Shift crew summaries

Reports

  1. Hall D Target Controls Restarted
  2. Follow-up Re: MAJOR alarms on solenoid cryogenics (Josh)
  3. Replaced fan tray of VXS crate D2-8-BOT DIRC
  4. FDC Wire Efficiency Update (Alex)
  5. New FCAL HVs (Mark)
  6. 2 FCAL base changed (Mark)
  7. New FCAL HVs (Mark)
  8. Numbers of triggers collected since start of production running
    • PARA 0/90 deg: 1.55B
    • PERP 0/90 deg: 2.46B
    • PARA 45/135 deg: 2.49B
    • PERP 45/135 deg: 2.43B
    • AMO: 1.08B

Problems

  1. Issue: Failed to retract amorphous radiator
  2. Program crushed after press analyze button when finishing Harp scan
  3. Runs won't start, Sergey excluded part of CDC
    1. HallD Run #131473: CDC occupancy plot

Plans

  1. Perform Harp scans
  2. Take 10 min raw-mode run
  3. Take production data
    • Start with remaining 121M events on amorphous radiator, then continue with diamond sequence


Thursday May 8

Last 24h

  1. Maintenance day, no beam since 06:30
  2. After maintenance: beam optics restore of accelerator
  3. Shift crew summaries

Reports

  1. FCAL bases failing (Mark)
  2. ECAL connector fixed (Sasha)
  3. Moller scattered beam monitors and scanner detectors installed (Sasha)
  4. Hall D Tagger Aluminum Dump Skid (Tim)
  5. Energy-dependence of rho decay asymmetry (Alex)
  6. Hall D LCW adjustment (Tim)
  7. Numbers of triggers collected since start of production running
    • PARA 0/90 deg: 1.55B
    • PERP 0/90 deg: 2.46B
    • PARA 45/135 deg: 2.49B
    • PERP 45/135 deg: 2.43B
    • AMO: 1.08B

Problems

  1. FCAL base (-13:-18) lockup unable to be cleared by resetting HV setpoints
  2. DIRC crate not reachable via network (Sergey)
  3. Issue of Alarm of TGTCRIO:hb_check from CONTROLS->LabView [26] (Alexandre)

Plans

  1. Beam not expected before 12pm
    • MCC will start with Hall A
  2. When beam is back: perform Harp scans
  3. Take production data
    • Start with remaining 121M events on amorphous radiator, then continue with diamond sequence